he arithmetic mean, or ordinary average, of a set of fifty numbers is 32. The average of a second set of seventy numbers is 53. Find the average of the numbers in the sets combined.
A) 44.25 B) 60 C) 42.5 D) 40 E) 50.25

Answer by edjones(8007)

50*32=1600
70*53=3710
.
1600+3710=5310
50+70=120
.
5310/120=44.25
.
Ed